LBC Cryostat

LBC Hub Focus Adjustment

  • Photo of position-adjusting bolts (loosen bolts, but do NOT remove them.
  • Current LBC Collimation models express mirror position, Z [mm], as a function of telescope elevation (deg) and upper C-ring steel temperature, TTEMP301 (deg C), using the following polynomials, where SX_Zgo and DX_Zgo are the global offsets [mm] for SX and DX sides, respectively.
    • SX: Z = SX_Zgo + 2.292 + -0.02179*elev + -0.0001865*elev*elev + 0.085*TTEMP301
    • DX: Z = DX_Zgo + 2.800 + -0.02476*elev + -0.0001788*elev*elev + 0.085*TTEMP301
    • Summer 2014: SX_Zgo = DX_Zgo = -1.3 mm. For the winter position, the hubs are moved 1mm away from M1 and the collimation models are edited to increase SX_Zgo and DX_Zgo by 1mm.
    • Plots for the current SX(blue) and DX(red) models for a range of temperature (solid curve TTEMP301 = 0 deg C; dashed curves from colder bottom to hotter at top).
      *collimtemp_summer.jpg
      models in use at this time (31Oct2014), which have SX_Zgo=DX_Zgo = -1.3 mm. MIRRORZ plotted for LBC data taken from 06-28 October 2014 when TTEMP301 varied between~2.8 and 12 deg C.
      collimtemp_summer.jpg
      *collimtemp_sum.jpg
      models for summer which are predicted to give equal temperature ranges: 0C to +15C; for both SX and DX. SX_Zgo= -0.9mm, DX_Zgo= -1.3mm. Data are the same as plotted in collimtemp_summer.jpg
      collimtemp_sum.jpg
      • collimtemp_win.jpg: models for winter which are predicted to give equal temperature ranges: -10C to +5C; for both SX and DX. SX_Zgo= +0.1mm, DX_Zgo= -0.3mm
        collimtemp_win.jpg
      • collimtemp_aut.jpg: an intermediate position predicted to give a temperature range from -5C to +10C. SX_Zgo= -0.4mm, DX_Zgo= -0.8mm
        collimtemp_aut.jpg

  • LBC Hub moves to Summer/Winter positions:
    • 08 May 2010 to Summer position
    • 01 Dec 2010 to Winter position (+1mm away from primary), TShih
    • 12 Nov 2011 +1mm Doug
    • 21 Apr 2012 to Summer position (1mm towards primary) J Morris & M Wagner
    • 14 Jan 2013 to Winter position (1 mm away from primary, red was reindexed) J Morris
    • 08 March 2013 caliper scales installed on hubs, readings: 85.00mm Blue, 94.00mm Red T Shih
    • 15 March 2013 to Summer position (1 mm towards primary, scale positions 84.00mm Blue, 95.08mm Red) J Morris
    • 17 March 2013 towards Winter position (0.5 Blue, 0.7mm Red away from primary, scale positions 84.52mm Blue, 94.40mm Red) J Morris
    • 27 April 2013 installed new brackets on caliper scales, new scale positions 84.62mm Blue, 94.12mm Red) J Morris
    • 25 May 2013. LBC hubs moved to summer position by E Solheid. Blue was moved 0.5mm towards primary and Red 0.75mm towards primary.
    • 10 July 2013 LBC hub scale readings: 84.12mm Blue, 94.72 Red J Morris
    • 16 Nov 2013 LBC-Red hub moved in X and Y, and collimation table changed accordingly:
      • Changed DX Xoffset from -1.614 to +0.386 after moving hub +2.0 mm in X (loosen 1/4 turn) to improve mixed copointing JMH 20131116
      • Changed Yoffset from -0.826 to -0.126 mm in the same adjustment. JMH 20131116
    • 26 Nov 2013 LBC hubs both moved to winter position (1mm away from primaries)
      • SX Zglobal set to -0.300 on 20131126 for winter hub position JMH
      • DX Zglobal to -0.3 for hub moved to winter position 20131126 JMH
      • Measurement on 29 May of this winter's position before move yields: Blue = 85.12mm Red = 94.2mm
    • 29 May 2014 LBC hubs moved to summer position (1mm toward primaries) Blue is now 84.11mm and Red is now 95.12mm
    • 18 Nov 2014 LBC hubs moved to autumn position (intermediate distance away from primaries, Blue -0.9mm, Red -0.5mm)
      • New positions: Blue = 84.92mm, Red = 94.60mm
    • 23 Dec 2014 LBC hubs moved to winter position (full distance away from primaries, Blue -0.5mm, Red -0.5mm from autumn position)
      • New positions: Blue = 85.42mm, Red = 94.12mm
    • 06 April 2015: LBC hubs moved to autumn/spring position. Both hubs were moved 0.5mm closer to M1.
      • New vernier positions: Blue = 84.92 and Red = 94.32 but Elliott notes that the scale on Red seems unreliable.
    • 13 June 2015: LBC hubs moved to summer position (0.5mm closer to M1 from autumn/spring position), performed by J Morris.
      • New vernier positions: Blue = 84.32 and Red = 94.60. At this time we are not confident that the vernier scales are changing accurately due to some binding of the mechanism, the 0.5mm move is measured by a digital indicator.
    • 13 Sep (June 2015): after the SANS crash the SX and DC PMLBCCollimation.dat files recovered were those from April. The summer hub values have been restored to the files now.
    • 30 Sep 2015 - set SX Zglobal to -1.40 by JMH
    • 03 Nov 2015: LBC Hubs moved to Autumn position by J Power.
      • LBCB Moved 0.59 mm away from M1, moving from 84.32mm on the LBC Blue position caliper to 84.91mm.
      • LBCR position caliper unreliable. Reading 94.62mm to start but looks snagged. Used 2 indicators with magnetic bases placed at different locations to ensure LBCR moved ok. Moved LBCR 0.511mm (1st measurement) 0.506mm (2nd measurement) away from M1. Position caliper reading 94.30 after.
    • 29 Dec 2015: LBC Hubs move to winter position by E. Solheid. Both moved 0.50 mm away from M1. New dial indicators installed on both cameras, will update procedure with photos.
    • 16 Feb 2016: LBC Hubs move by J. Power. LBC blue moved 0.542mm closer to M1 (spring/autumn position); LBC red moved so net is spring/autumn position value now, but it was moved 0.497mm away from M1 to account for incorrect physical position of the Hub before. SX (LBC-Blue) is now -0.84 in config file, and DX (LBC-Red) is now -0.8 in config file.
    • 06 Jun 2016: LBC Hubs moved by A. Cavallaro - both hubs moved 0.5mm closer to M1 for summer position. SX = -1.34, DX = -1.30
    • 30 Nov 2016: LBC Hubs moved to full winter position by E. Solheid
    • 16 Apr 2017: LBC Hubs moved to autumn/spring position by A. Cavallaro.
    • 09 June 2017: LBC Hubs moved to summer position by A. Cavallaro (SX at -1.34 and DX at -1.30)
    • 23 Sept 2017: LBC Hubs moved to spring/fall position by E. Solheid
    • 08 Dec 2017: LBC Hubs moved to winter position.
    • 10 Apr 2018: LBC Hubs moved to spring/fall position by E. Solheid
    • 8 May 2018: LBC Hubs moved to summer position by J. Morris.
      • This time, after noticing that the SX Z positions lay ~0.3mm above the collimation model, a move of -0.7mm was requested in SX, instead of the usual -0.5mm. It turned out to be not possible to resolve moves on such a fine scale. John estimated he moved SX by -0.53mm and DX by -0.5mm DX.
      • Also, having similar global offsets led to widely different starting pupil sizes, because the constant in the 2nd order polynomial fit to mirror Z and elevation was different for SX and DX. So both the Z global offsets and Offset Collimation Coefficients were edited for SX and DX. The following summarizes the changes made first to equalize the pupil sizes and then to accomodate the hub moves:
    • 5 Oct 2018: LBC Hubs moved to fall position, Geno and Elliott.
    • 8 Dec 2018: LBC hubs moved to winter position by G. Bechetti.
    • 23 April 2019 LBC hubs moved to spring position by Elliott
    • 22 June 2019 LBC hubs moved to summer position by Elliott
    • 6 October 2019 LBC hubs moved to fall position by Elliott
    • 30 Nov 2019 LBC hubs moved to winter position by Chris & Elliott
    • 30 May 2020 LBC hubs moved to summer position by Jared and Rick H
    • 14 Oct 2020 LBC hubs moved to fall/spring position by Jared and Elliott
    • 16 Dec 2020 LBC hubs moved to winter position by Elliott
SX_Z_Global SX_Offset DX_Z_Global DX_Z_Offset
-0.84 2.292 -0.80 2.800 prior to 20180508 (S/DXPMLBCCollimation_20180508.dat)
-0.34 2.292 -1.00 2.800 20180508 - to bring pupil diameters closer to the expected 39pix for dZ=-0.5mm (*_20180509.dat)
-1.00 2.422 -1.00 2.300 20180509 - after hubs moved -0.53mm(SX) and -0.50mm(DX) to summer positions ("live" versions)
Updating the Collimation Tables for SX and DX:
  • The next step after the hub has been moved is to update the collimation tables for SX and DX.
  • Login to obs1 with the "telescope" account, and from there login to tcs1 as user tcs.
  • cd /lbt/data/config/tcs/PSF
  • the two files to edit are: DXPMLBCCollimation.dat and SXPMLBCCollimation.dat
  • Using your favorite editor (i.e. emacs) open each file.
  • Make the changes to the "dz" column (1st row) in the Global Offsets Section (in the "Elevation Correction Coefficients Section, see attached jpg, highighted in red) for SX
  • collimation_table.jpg
  • Please update the time and amount changed in the header comments section (noted by red arrow in above picture)
  • Repeat this for DX
  • The tables will be ingested and update once LBC is authorized (or re-authorized).

LBC Filter Wheel Motor Current History

The following table represents the GRC (get real current) value (milliamps) from the LBC filter wheel motors during the balance testing -- maximum and average values. This is for historical purposes. We run the balance tests with the telescope at zenith or horizon.

  Blue1 Blue2 Red1 Red2
Date Avg Max Avg Max Avg Max Avg Max
May-2015         284 zenith
273 horiz
350 zenith
320 horiz
244 zenith
243 horiz
280 zenith
290 horiz
Oct-2014 370 zenith
394 horiz
580 zenith
680 horiz
268 zenith
389 horiz
580 zenith
790 horiz
       
Sep-2014 413 horiz 700 horiz 390 horiz 790 horiz        
Aug-2014 763 zenith
761 horiz
950 zenith
1010 horiz
302 zenith
225 horiz
580 zenith
600 horiz
237 zenith
204 horiz
270 zenith
230 horiz
161 zenith
159 horiz
200 zenith
190 horiz
Aug-2013 288 530 325 680 411 480 156 190

LBC History

  • 06 June 2010 LBC Blue goes soft*
  • 10 June 2010 Some problems with small turbopump
  • 11 June 2010 New vacuum on LBCBlue. LN2 filled

Warm ambient temperatures! Controller overheating
  • 15 June 2010 2 New .21A fans replace 3 .13A fans

Summer Shutdown 2010
  • 21 June 2010 Last LN2 Fill
  • 4 July 2010 Computers out of treehouse
  • 6 July 2010 Electronics and wiring out of treehouse

Fall Startup 2010
  • 30 August 2010 Electronics go back in treehouse
  • 8 September Cryostats installed. Finish final cabling.
  • 9 September Wiring connecting Blue camera Profibus rotator encoder discovered damaged and repaired.
  • 17 September Broken wire discovered in video cable for chip 2, Red camera. Dave Baxter at ITL repairs it.

Fall Operations
  • 8 October 2010 Red camera filter wheel #1 reported stuck, not operating correctly.
  • 6 November setscrews attaching red filter wheel #1 encoder to shaft were installed by Fernando.
  • 8 November Vacuum in Red camera noticed increasing from 10-5 to 2.8x10-4
  • 17 December Red camera removed from telescope for pumping in lab.
  • 25 January 2011 Installed new KVM switch in treehouse
  • 17 March 2011 Turn off Blue+HK and decable. J-box removal/reroute cables for ARGOS swingarm.

Summer 2011
  • 12 Aug 2011 Filters of both cameras cleaned.

Spring 2012
  • LBC Red Detector Issues
  • 15 June 2012 Red tech controller swapped
  • 16 June 2012 #2 fiber in barrel/rotator changed to #4 on Red tech controller

Spring 2014
  • It had been found that the L1 baffle on LBC-Blue was not on. Don't know for how long it was missing.
  • Elliott found it and reinstalled it on 20140529.

Spring 2017 (resuming this log...)
  • LBCB with controller #0 started to show banding on biases & images (IT 6452).
    • Note that controller #0 has a thermal problem and needs 2 fans which James & Jenny had installed in it.
    • Contoller #3 was swapped in, but found to show vertical striping and moderate light levels, although not on biases (IT 6549).
    • Controller #6 was found and swapped in.
    • Controller #0 shipped back to SkyTech and quotation for repair of #3 was requested from SkyTech on 04/21 (RR email 04/25/2017).
  • LBCB with controller #6:
    • Kellee updated conf/bluechannel.conf file to set CCDHardwareID to 6.
    • Kellee adjusted the voltages in hwmodel_C/bsci/normal/CDS_A600_Offset.ini to aim for bias levels of ~500 adu on all 4 chips, and she updated the lbc.zip file.
      • We used Fernando's guidance: a voltage offset 0.02 corresponds to a change of -250 adu in the bias level to estimate the new voltage settings, and this worked well.
    • a set of data was taken to measure gain & linearity.
      • Calib_OBs/ClosedDomeTesting/gi2BR_dome.ob will take a pair of exposures through g(LBCB) and i(LBCR). With the lower floods, you get about 24k and 7k counts in blue and red, respectively. Exposure times started at 0.3 sec and went up to 3.5 sec, however the last datapoints with counts >50k were not useful (although the detector appears linear, signal-variance started to turn over).
      • data were taken with LBCs out of the beam, where light incident at high angles causes some interference & scattering of optical surfaces. This would elevate the variance.
      • signal-variance method was used, where the signal represents the average of the midpts returned by imstat (unintentionally not clipped) over a 50x50 region for an image pair that was taken with the same exposure time. The variance is the stddev measured by imstat over the same 50x50 region of the difference between the two images in a pair, where one has been scaled, by the ratio of their midpts, to match the other. On the plot of signal vs variance, the slope of the best-fit line represents the gain. Linear fits were made using all points by the last 4, where counts were ~50,000 or higher, and the resulting gains, for 6 points on each chip, are given in the table below:

LBCB gains (Controller #6)
region gain_1 gain_2 gain_3 gain_4
rotcen50 1.54 1.76 1.69 1.79
topleft50 1.74 1.33 1.71 1.57
topright50 1.25 1.82 1.55 1.43
center50 1.50 1.26 1.34 1.38
bottomleft50 1.43 1.62 1.75 1.63
bottomright50 1.26 1.63 1.75 1.67

  • LBCR with controller #1. Using a similar analysis on the 50x50 region about the rotator center, gains for LBCR were found to be:

LBCR gains (Controller #1)
region gain_1 gain_2 gain_3 gain_4
rotcen50 1.84 1.97 1.96 1.95
September 2017

We resumed from summer shutdown on September 1/2, using LBCB (controller #6) and LBCR (controller #1). We noted that LBCR has been showing some intermittent low level (1-2% of the background or 5-10 ADU) banding in biases and low background (<~600 ADU) observations; at higher background levels more typical of science exposures, I think it is present but just swamped by the sky. We have been actively troubleshooting this, examining temperature effects, noise from other electronics/cables and the power supply. All of the troubleshooting efforts made are detailed in IT #6748. On the weekend of September 9/10, James Riedl swapped in the controller #0 that had been recently serviced by Skytech. We had some fiber issues with this controller, still saw intermittent banding and so he swapped back in controller #1. He tried a new power supply which had lower noise and, though it could not be rigged up to be used right away, plans will be made to swap out the aging power supplies. He separated the cables carrying the clock voltages from those carrying bias and video. Still further tests showed intermittent banding. The banding extends into the overscan and can be subtracted out in the reduction, using not the polynomial fitting algorithms but instead ccdproc.function = (mean, median or minmax).

One thing I did learn in the troubleshooting is that, when the filter is moved during readout, the image read out is often banded, and until resolved, this argues against setting up an OB which cycles through filters at a given dither position.

20 Sep 2017 Controller #3 received back from Skytech.

Feb 2020 LBCR trackers not coming up due to a bad input voltage (IT 8048). The controller (#3) was swapped out and #0 installed in its place. With controller #0, the red trackers came up, but the images showed horizontal banding and we also had some problems with another bad input voltage. Work was done on controller #3 in the lab and it was swapped back in. Red trackers with refurbished #3 work well.

LBC Filter Cleaning

This page was started in 2016B to list the dates on which LBC skyflats were taken, together with any changes to the instrument, such as condensation and subsequent cleaning of a filter or lens, which would render LBC skyflats taken before unsuitable for use with data taken afterwards. It has been moved from the old sciops cookbook (wordpress) to the Twiki for now and will be updated here. Note that after November 2016, dates of skyflats have not been recorded, but I think the cleanings have been.

Date LBCB Filters LBCR Filters Action (Sky Flats or other)
20160905 Uspec,B-Bessel,V-Bessel V-Bessel, r-SLOAN, Y-FAN, z Sky Flats; only 2 good z flats obtained
20160919 V-Bessel, Uspec I-Bessel, i-SLOAN, z-SLOAN filters cleaned
20160925 B-Bessel, V-Bessel V-Bessel, R-Bessel Sky Flats; star residuals even after masking in blue flats
20160929   i-SLOAN? filters cleaned; (?=the filter most exposed when the wheels were parked)
20161007 g-SLOAN, r-SLOAN V-Bessel, r-SLOAN, i-SLOAN Sky Flats; some cirrus
20161011 B-Bessel, g-SLOAN, r-SLOAN V-Bessel, i-SLOAN, z-SLOAN Sky Flats; very bright star which could not be fully masked in r-SLOAN and z-SLOAN flats
20161029 U-Bessel, B-Bessel, V-Bessel V-Bessel, R-Bessel Sky Flats in morning twilight
20161029 all all filters cleaned; see note in IT #6231
20161119 U-Bessel, B-Bessel R-Bessel Sky Flats; R-Bessel flats show 4 circles from dust or drops (not on an I-Bessel test flat). IT #6231
20161119   R-Bessel R-Bessel cleaned
20170529   R-Bessel R-Bessel cleaned after debris from moth (IT #6627)
During the summer 2017, all filters were cleaned in Tucson and LBCB U-Bessell,B-Bessell,V-Bessell,Uspec and LBCR F972N20 were scanned at NOAO
Also, LBCB and LBCR lenses were cleaned on 20170903(day) and 20170725, respectively.
 
20170902 Uspec, V-Bessell z, V-Bessell Sky Flats
20170903 B-Bessell I-Bessell Sky Flats
20170903(day)     all LBCR filters cleaned (See pictures. Labels {2,4,5,2',3'} correspond to {r,i,z,R,V})
20170905 Uspec V-Bessell Sky Flats (telescope not tracking)
20170906 Uspec, B-Bessell, r-SLOAN, g-SLOAN(only 2) r-SLOAN, i-SLOAN, V-Bessell, I-Bessell (only 2), F972N20 Sky Flats
20170907 U-Bessell, g-SLOAN V-Bessell, I-Bessell Sky Flats
20170910 B-Bessell, V-Bessell R-Bessell, z-SLOAN Sky Flats
20170913(day)     LBCR lenses L1 and L2 cleaned (smaller ones behind, L3-L5?, could not be accessed but look dirty)
20170918 V-Bessell, g-SLOAN TiO _784, Y-FAN Sky Flats
201710xx     Sky Flats
20171025     moths vacuumed out from behind L1 on both LBCR and LBCB
20180707     on LBCB J. Riedl cleaned both sides of L1 and L2 and the side of L3 facing L2
2018 August: LBCB and LBCR filters and L1 and L2 were all cleaned before restart on 20180901
2019 July/August: LBCB and LBCR filters and L1 and L2 were all cleaned before restart on 20190901
20190910 g-SLOAN   moth excrement cleaned off of g-SLOAN
20191203   i-SLOAN Two drops cleaned off of i-SLOAN (cb)
20191221   R-Bessell During the night (20191221 UT), a spot appeared on R-Bessell images (IT 7980). Morning twilight flats were obtained, and the filter was cleaned a few hours later. (cb)
20200123 all (U-,B-,V-Bessel, g-, r-Sloan) all (V-,R-,I-Bessel,r-,i-,z-Sloan,Y,F972N20) LBCB r-Sloan, LBCR F972N20 and i-SLOAN cleaned, all others dusted
2020 August: LBCB and LBCR lenses and filters were cleaned during summer shutdown. The outside of the dewar window of LBCR was cleaned. Debris was noted to be inside the window (IT 8130) and that was not cleaned.
20200927 g-SLOAN   A spot was cleaned after noticing pseudo-flats showed a smudge on chip 1 (cb)

LBC Flats

In 2017B, a log of LBC twilight sky flats was started. This tables at this link provide the numbers of "good" and "acceptable" sky flats that were obtained each month through each camera+filter combination. The table entries link to lists of the filenames and median count levels.

LBC Troubleshooting

See the LBC Troubleshooting page for an architecture diagram, hardware failure scenarios, link to the computer swap procedure, links to other troubleshooting notes.

  • Moths inside L1 on LBCB and LBCR 2017-Oct-16:
    LBCR_moths_L1_2017oct16.jpg

  • Moths inside L1 on LBCB and LBCR 2017-Oct-16:
    LBCB_moths_L1_2017oct16.jpg

LBC CCD Cameras Maintenance Log

Annual maintenance activities as requested in I602s00114 "Procedure For LBC Annual Preventive Maintenance" (Step 2).

Procedure: Lab Setup Instructions

Topic attachments
I Attachment Action Size Date Who Comment
20170902_2_r-SLOAN.jpgjpg 20170902_2_r-SLOAN.jpg manage 1 MB 10 Sep 2017 - 18:09 OlgaKuhn debris on LBCR filters noticed 20170902 and cleaned during the day on 3 September
20170902_2p_R-Bessell.jpgjpg 20170902_2p_R-Bessell.jpg manage 1 MB 10 Sep 2017 - 18:10 OlgaKuhn debris on LBCR filters noticed 20170902 and cleaned during the day on 3 September
20170902_3p_V-Bessell.jpgjpg 20170902_3p_V-Bessell.jpg manage 1 MB 10 Sep 2017 - 18:11 OlgaKuhn debris on LBCR filters noticed 20170902 and cleaned during the day on 3 September
20170902_4_z-SLOAN.jpgjpg 20170902_4_z-SLOAN.jpg manage 1 MB 10 Sep 2017 - 18:11 OlgaKuhn debris on LBCR filters noticed 20170902 and cleaned during the day on 3 September
20170902_5_i-SLOAN.jpgjpg 20170902_5_i-SLOAN.jpg manage 1 MB 10 Sep 2017 - 18:11 OlgaKuhn debris on LBCR filters noticed 20170902 and cleaned during the day on 3 September
2017BFlats_updated_31Jan.xlsxxlsx 2017BFlats_updated_31Jan.xlsx manage 47 K 01 Feb 2018 - 00:03 OlgaKuhn LBC twilight flats 2017B
DXPMLBCCollimation.jpgjpg DXPMLBCCollimation.jpg manage 311 K 19 Jan 2013 - 00:36 JohnHill Plot of DX Primary Mirror Collimation Model
LBCB_moths_L1_2017oct16.jpgjpg LBCB_moths_L1_2017oct16.jpg manage 1 MB 26 Oct 2017 - 00:17 OlgaKuhn Moths inside L1 on LBCB and LBCR 2017-Oct-16
LBCR_moths_L1_2017oct16.jpgjpg LBCR_moths_L1_2017oct16.jpg manage 1 MB 26 Oct 2017 - 00:17 OlgaKuhn Moths inside L1 on LBCB and LBCR 2017-Oct-16
LBC_Hub.pdfpdf LBC_Hub.pdf manage 851 K 06 Dec 2010 - 21:15 JohnMorris  
SXPMLBCCollimation.jpgjpg SXPMLBCCollimation.jpg manage 314 K 19 Jan 2013 - 00:34 JohnHill Plot of SX Primary Mirror Collimation Model
collimation_table.jpgjpg collimation_table.jpg manage 709 K 14 Jun 2015 - 03:44 BarryRothberg  
twiflats_BBb.datdat twiflats_BBb.dat manage 5 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_BS.datdat twiflats_BS.dat manage 4 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_BUb.datdat twiflats_BUb.dat manage 1 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_BVb.datdat twiflats_BVb.dat manage 3 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_BgS.datdat twiflats_BgS.dat manage 6 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_BrS.datdat twiflats_BrS.dat manage 6 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_RC.datdat twiflats_RC.dat manage 562 bytes 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_RF.datdat twiflats_RF.dat manage 1 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_RIb.datdat twiflats_RIb.dat manage 3 K 11 May 2018 - 05:47 OlgaKuhn  
twiflats_RRb.datdat twiflats_RRb.dat manage 4 K 11 May 2018 - 05:48 OlgaKuhn  
twiflats_RT.datdat twiflats_RT.dat manage 1 K 11 May 2018 - 05:49 OlgaKuhn  
twiflats_RVb.datdat twiflats_RVb.dat manage 5 K 11 May 2018 - 05:49 OlgaKuhn  
twiflats_RY.datdat twiflats_RY.dat manage 1 K 11 May 2018 - 05:45 OlgaKuhn  
twiflats_RiS.datdat twiflats_RiS.dat manage 3 K 11 May 2018 - 05:37 OlgaKuhn lists of twilight flats from 2017 September to 2018 May 11 UT
twiflats_RrS.datdat twiflats_RrS.dat manage 4 K 11 May 2018 - 05:47 OlgaKuhn  
twiflats_RzS.datdat twiflats_RzS.dat manage 3 K 11 May 2018 - 05:46 OlgaKuhn  
Topic revision: r96 - 16 Dec 2020, ElliottSolheid
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback